d) Consider a perfectly competitive market with a market demand curve that is given by the equation P 2000-2. A representative firm in this market has a total cost curve given by the equation TC 121 64q+ and a marginal cost curve given by MC- 642g. O is the market quantity and q is the firm quantity. Suppose the short-run price (P) in this market is S100 i) What is the market quantity in this market given this short-run price? (2 marks) ii) What is the representative firms level of production given this short-run price? (3 marks)iii) What is the representative firms level of profits in the short-run given this market price? (3 marks) iv) Can this short-run equilibrium also represent a long-run equilibrium for this firm? Explain your answer. (2 marks) v) Rounding to the nearest whole number, how many firms are operating in the short-run in this market given a market price of 100 firms? (2 marks)

Please explain in details with step-by-step solution, Thank you very much
Answer #1

Ans (i) The Demand curve of the firm is given as; P= 2000-Q

Total Cost Curve is given as ; TC = 121+ 64q + q2

Marginal Cost curve is given as ; MC = 64 + 2q

The market quantity can be calculated as follows:-

Since P = $100 , putting this value in the market demand curve we get ,

P= 2000-Q

$100 = 2000-Q , so Q = 1900 which is the market quantity.

Ans (ii) We know that for a firm operating in a competitive environment, the marginal revenue received is always equal to the market price. Therefore, a firm operating under competition will always produce at the level of output where the Marginal cost of the last unit produced is just equal to the market price.

So, MC = Market Price

64+2q=100

Solving this we get q = 18 which is the representative firm's level of production.

Ans (iii) The representative firm's level of profit occurs at a point where MR=MC.

Now, we can say that the MR curve is half of the market demand.

So, Marginal Revenue = i/2 *1900=950

Now, We know that Profit maximization will be at the point where MR=MC

So, 950=64+2q

Solving this we get q = 443 which is the firm's profit.

Ans (iv) In microeconomics, the long run is the conceptual time period in which there are no fixed factors of production so that there are no constraints preventing changing the output level by changing the capital stock or by entering or leaving an industry. The long-run contrasts with the short run, in which some factors are variable and others are fixed, constraining entry or exit from an industry. In macroeconomics, the long run is the period when the general price level, contractual wage rates, and expectations adjust fully to the state of the economy, in contrast to the short run when these variables may not fully adjust.

In the long run, firm change production levels in response to economic profits or losses and the land, labour, capital goods and entrepreneurship vary to reach the minimum level of long-run average cost. The long run is associated with the long run average cost (LRAC) curve in macroeconomic models along which a firm would minimize its average cost (cost per unit) for each respective long-run quantity of output.

The transition from the short run to the long run may be done by considering some short-run equilibrium that is also the long -run equilibrium as to supply and demand, then comparing that state against a new short-run and long-run equilibrium state from a change that disturbs the equilibrium.

Ans (v) The number of firms in the short-run can be calculated as follows:-

We have calculated the market quantity in part (i) as 1900. Now the market price is given for 100 firms. By dividing the market quantity by 100 we get the number of firms =1900/100 =19.